i just had this question answered but i think there is more to it, can someone please help me. the problem is (-10xy^3z^2)^3 the answer is -1000x^3y^9z^6 but my question is don't you have to remove the negative 1000 by changing the sign and flipping it? to make the answer x^3y^9z^6/1000? thanks so much in advance for answering!
Answer by Nate(3500)

1000^(-1) = 1/1000
-1000 = -1000
(-10xy^3z^2)^3 = (-10)^3(x)^3(y^3)^3(z^2)^3 = (-10)(-10)(-10)(x)(x)(x)(y^3)(y^3)(y^3)(z^2)(z^2)(z^2) = -1000x^3y^9z^6
If that does not help, contact me at codestumper@hotmail.com .